OpenStudy (anonymous):

solution to 3x-3y=12 and 4y-8x=-16

OpenStudy (anonymous):

3x - 3y = 12 4y - 8x = -16 12x - 12y = 48 -24x+12y = -48 -12x = 0 x = 0 3(0) - 3y = 12 -3y = 12 y = -4